Why must chemical equations be balanced?

In a balanced chemical equation, the coefficients in front of each reactant and product allow us to predict whether the given reaction is possible or not.
In a balanced chemical equation, the coefficients in front of each reactant and product allow us to predict whether the reaction requires a catalyst or not.
In a balanced chemical equation, the coefficients in front of each reactant and product allow us to predict the amount of product that can form from a given amount of reactant.
In a balanced chemical equation, the coefficients in front of each reactant and product allow us to predict how much of one reactant is required to react with a given amount of another.
In a balanced chemical equation, the coefficients in front of each reactant and product allow us to predict how much reactants are required to form a given amount of products.
Answer #1

consider a chemical equation

aA + bB ---> cC + dD

here

a, b,c and d are the coefficients

from the above equation it is clear that

a moles of reactant A on reaction gives c moles of product C

moles of product C formed = ( c /a ) x moles of reactant A taken

so

In a balanced chemical equation, the coefficients in front of each reactant and product allow us to predict the amount of product that can form from a given amount of reactant.

a moles of reactant A reacts with b moles of reactant B

moles of reactant A required = ( a / b) x moles of reactant B taken

so

In a balanced chemical equation, the coefficients in front of each reactant and product allow us to predict how much of one reactant is required to react with a given amount of another.

also

from the reaction

to get c moles of product C , we require a moles of reactant A

moles of reactant A required = ( a/c) x moles of product C formed

so

In a balanced chemical equation, the coefficients in front of each reactant and product allow us to predict how much reactants are required to form a given amount of products.